OpenStudy (anonymous):

solve n check 5+2x-1=7x +10-3x

OpenStudy (owlcoffee):

Let's first recognize the problem we hace in front of us, we only have one variable and it's has an exponent of one (don't worry if you don't know this yet). So we say this is a "single variable equation" or "1st degree equation". The way to solve it is leaving the numbers with the variable on one side (the numbers with an "x" on them) of the equal sign, nd the constants (numbers) on the other: \[5+2x-1=7x+10-3x\] So, let's add 3x on both sides, and sustract 7x too: \[3x-7x+5+2x-1=10\] Let's now sustract 5 and add 1: \[3x-7x+2x=10-5+1\] Operate: \[3x-7x+2x=6\] Now, to add up the numbers with variables, we enclose them in a parenthesis and leave the x outside: \[(3-7+2)x=6\] And do this: \[x=\frac{ 6 }{ (3-7+2) }\] \[x=\frac{ 6 }{ 2 }\] \[x=3\] Nw, replace all the x's for 3's, and see what happens.